OpenStudy (hocuspucus292333393):

Help Evaluate 4m+23+np−3, when m=7, n=2, and p=8. Enter your answer in the box.

jimthompson5910 (jim_thompson5910):

Replace m with 7 Replace n with 2 Replace p with 8 Then evaluate using PEMDAS \[\Large 4m+23+np-3 = 4*7+23+2*8-3 = ???\]